OpenStudy (anonymous):

An electron in a cathode-ray tube is traveling horizontally at 2.16 x10^9 cm/s when deflection plates give it an upward acceleration of 5.43x 10^17 cm/s2. a)(a) How long does it take for the electron to cover a horizontal distance of 6.07 cm? (b) What is its vertical displacement during this time? Can some please help me understand how to even start this problem off? I don't understand how to use the acceleration in this? Do I treat it as a normal vector?

OpenStudy (anonymous):

It isn't quite as complicated as the question makes it sound. We just need to identify the pieces of information. Intiial conditions: v(horiz) = 2.16x10^9 cm/s v (vertical) = 0 a (vertical) = 5.43x10 cm/s^2 Part a can be solved with d= v(initial)*t (use the horizontal components) Part b can be solved with d= at^2 (use the vertical components) Both of the above eq'ns are part of the displacement eq'n: d= d(initial) + v(initial) + at^2 You just are applying them to a novel situation, but the procedure is the same as if you are solving a normal projectile problem